OpenStudy (anonymous):

What is the standard form of the number in the calculator display? 2.3e-9

OpenStudy (adajiamcneal):

Reorder the terms: -9 + 3e = 0 Solving -9 + 3e = 0 Solving for variable 'e'. Move all terms containing e to the left, all other terms to the right. Add '9' to each side of the equation. -9 + 9 + 3e = 0 + 9 Combine like terms: -9 + 9 = 0 0 + 3e = 0 + 9 3e = 0 + 9 Combine like terms: 0 + 9 = 9 3e = 9 Divide each side by '3'. e = 3 Simplifying e = 3

OpenStudy (anonymous):

e stands for too many zeros on a calculator like when its 2.3 * 10 to the power of something

OpenStudy (adajiamcneal):

oh 0.0000000023

OpenStudy (anonymous):

thx

OpenStudy (adajiamcneal):

welcome :)